Show that f(z) = x-iy and g(z) = Im(z) do not satisfy the Cauchy-Riemann equations.
Answer #1

f(3) = x-iy J u= 0 ay-o local v=-4 Vocao, Vy=-1 4x + Vy Myt -voc . f(z) does not satisfy CR equations.z=xtiy g(z) = in (z) 9(2) = iy u=4 =o Cu / Vocco, Vy=0 Ux= vy but 9(z) does not CR equation. Uy t-Voc - =0+=Uu satisfy